# On the fate of spacetime singularities

### v1

@@ Abstract
I investigate spacetime singularities from the point of view of the wavefunction of the universe. In order to extend the classical notion of geodesic incompleteness one has to include the proper time of an observer as a degree of freedom in the Wheeler DeWitt equation. This leads to a Schrödinger equation along the observer worldline. Near the singularity, as in the classical BLK treatment, I ignore spatial gradients and effectively describe the spacetime around the worldline in the mini-superspace approximation. The problem proves identical to a spherically symmetric scattering of a quantum particle off a central potential and singularity avoidance is tantamount to the unitary evolution of this system. Standard types of matter (dust, radiation) correspond to regular potentials and thus lead to a bounce. The most singular component, spatial anisotropy, is associated to a conserved charge and yields a negative inverse-square potential—like standard angular momentum, but with opposite sign. This potential is critical, in that the unitarity of the evolution depends on the actual numerical factor in front of it, i.e., to the anisotropy charge.
@@ I Introduction
However, this conclusion might be too pessimistic. As long as the gravitational wavefunction maintains its support on smooth sub-Planckian metrics it should be under the control of the low-energy theory, but can still display features that cannot be reproduced by a point-like classical system. For example, quantum tunneling is made possible by the non-vanishing width of the wavefunction, which allows it to “feel” regions of the potential away from its peak. In this case, a weak-coupling semiclassical treatment is indeed available, but it applies to the Euclidean (rather than Lorentzian-) path integral, so it is not directly associated to a classical behavior. In gravity, semiclassical techniques apparently oblivious to UV physics have greatly contributed to the understanding of black hole thermodynamics (e.g. Gibbons and Hawking (1977); Hawking and Page (1983); Almheiri et al. (2021) to cite a few).
@@
However, as emphasized above, it should be the proper time of an observer, not some arbitrarily chosen field, that diagnoses a singularity. As the WdW equation only deals with dynamical fields we should introduce the trajectory of the observer and its proper time $\tt{t}$} as new degrees of freedom. While we send the observer into the singularity we can compute $\Psi(q^{a},{\tt{t}})$ along the worldline and obtain a probability conditional to the actual proper time $\tt{t}$} of the observer. Equivalently, we can compute the expectation values of the $q^{a}$ variables at given $\tt{t}$}.
This seems the correct starting point to extend the notion of singularity—or avoidance thereof—into the quantum regime. Without incorporating the clock variable $\tt{t}$}, inspection of the “standard” WdW wavefunction $\Psi(q^{a})$ generally leads to inconclusive results.
To clarify this, let me consider a very stupid example with only two fields, the scale factor $a$ and a scalar $\phi$. Say that $\Psi(a,\phi)$ turns out to be peaked along the field space trajectory $\phi=a$. The relational sense of this is clear Piazza and Vareilles (2025b). If, e.g., we use $a$ as our clock, we have, roughly, $\langle\phi\rangle=a$. Does this model contain a singularity or a bounce?
Only when $\tt{t}$} is added as a degree of freedom can we address this question. From $\Psi(a,\phi,{\tt{t}})$, one can compute the average values of $a$ and $\phi$ as functions of $\tt{t}$}. The probe should only slightly perturb the system, so $\Psi$ is still be peaked on a trajectory belonging to the two-dimensional plane $a=\phi$. However, this can include very disparate situations. We could have, say, $\langle a\rangle=\langle\phi\rangle={\tt{t}}^{2}+1$, which clearly represents a bounce. Or, we could find that both $\langle a\rangle$ and $\langle\phi\rangle$ tend to infinity at ${\tt{t}}=0$—more precisely, that $\Psi$ does not evolve unitarily in $\tt{t}$}. In this case, we would conclude that the low-energy theory alone is not able to resolve the singularity.
@@ II The Schrödinger equation
Here I review a simple way to include the proper time of an observer into the WdW formalism (see also Teitelboim (1983); Unruh and Wald (1989); Peter (2018); Malkiewicz et al. (2022); Burns et al. (2022); Kaya (2023); Bergeron et al. (2025)). By following Piazza and Vareilles (2025a), I will argue that the inclusion of an observer’s clock into the system turns the WdW equation into a Schrödinger equation along the worldline of the observer.
Quid est Tempus? — Classical GR offers a clear unambiguous answer to this question, at least locally, i.e. along a given observer’s worldline. The invariant line element, or proper time, $d{\tt{t}}$ is obtained by contracting the metric tensor $g_{\mu\nu}$ along the observer’s trajectory—see eq. (2) below. The actual operational procedures that the observer needs to implement in order to measure proper time are not of concern here. By definition, a good clock must tick homogeneously in the $\tt{t}$} variable.
@@
Quantum mechanically, the variance associated with $\tt{t}$} can be interpreted as the uncertainty to set the origin (e.g. midnight) of our clock variable with respect to all other fields of the theory along the worldline. Or, as Witten nicely puts it Witten (2024),
@@
This gauge corresponds to fixing the spatial coordinates $x^{i}$ in such a way that the shifts $N_{i}=0$. This can be done by identifying a congruence of geodesic observers, among which we can set at $x^{i}=0$ the specific observer that we provide with the clock (2). The formal way to do so is to introduce a non-relativistic fluid type of matter, described by three scalar fields $\phi^{I}$ where $\phi^{I}=constant$ represent the trajectories of the volume elements of the fluid, or of the observers inside the congruence. By adopting the unitary gauge choice $\phi^{I}=x^{i}$, the momentum constrain then reduces to $N_{i}=0$ (e.g. Nicolis et al. (2022), App. D). The Hamiltonian constraint is the WdW equation, that I discuss shortly. When both $N$ and $h_{ij}$ are functions of $t$ and $x^{i}$, the metric (4) is completely general.
@@
where $N{\cal H}$ is the Hamiltonian of the system without the clock. In the standard WdW approach one poses ${\cal H}\Psi(q^{a})=0$ ($q^{a}$ are now zero-dimensional mini-superspace quantum variables). By including the clock into the system, the WdW equation becomes ${\cal H}_{\rm tot}\Psi(q^{a},{\tt{t}})=0$. Upon use of the canonical relation $p_{\tt{t}}=-\,i\partial_{\tt{t}}$ this is just the Schrödinger equation as advertized,
@@ III Approaching the singularity
A convenient parameterization — In the absence of anisotropic stresses, as shown in some detail in Nicolis et al. (2022), the metric $h_{ij}$ can be conveniently parametrized as
@@
where $\Delta^{\!(2)}=\partial_{+}^{2}+\partial_{-}^{2}$ is the Laplacian in the flat two-dimensional anisotropy space and, as opposed to the schematic (7), all factors of $\alpha$ and $L$ have been made explicit. By measuring time in units of $L$, we can effectively set $L=1$ from now on.
@@
As predicted, anisotropies contribute a negative effective potential and act as a centripetal force. This is the analogous to having $\ell(\ell+1)<0$ in the Hydrogen atom. I now review what is the maximally singular behavior for the potential to be compatible with unitary evolution. I will show that typical matter ($w=0$ for non-relativistic matter, $w=1/3$ for radiation) contributes to the potential in a regular way and that the only possible source of instability is given by anisotropy, depending on its actual amount, i.e. on the value of $Q$.
@@ IV Unitarity
with $\gamma\geq 0$ and $\beta\geq 0$. Classically, this system is unstable because the Hamiltonian is unbounded from below. This allows potential energy to be converted indefinitely into kinetic energy, leading to runaway singular solutions, as in the case of a radial infall toward the origin. Quantum mechanically, however, the system (23) is regular if Landau and Lifshits (1991) (see also Perelomov and Popov (1970); Coon and Holstein (2002); Bawin and Coon (2003))
@@
The RHS of the above inequality is itself a function bounded from below if $\beta<2$, or if $\gamma<1/4$ in the marginal case $\beta=2$. Remarkably, this simple heuristic argument reproduces the correct power-law behavior of the potential and, in $d=3$, also the critical numerical value of the coefficient as in (24). According to the classical analysis of Landau and Lifshits (1991) (Sec. 35) a potential more singular than (24) simply leads to the infall of the quantum particle into the center.
@@
When $\beta>2$ the second term on the LHS of (28) dominates, and it cannot be set to zero by any obvious choice of $\lambda$. What happens in this case is that the phase of the wavefunction increases indefinitely while approaching the singularity. This can be seen Perelomov and Popov (1970) already at leading order in the semi-classical expansion $R=e^{i(S_{0}/\hbar+S_{1}+\dots)}$, which gives
@@ V Discussion
Classically, the energy density associated with spatial anisotropy scales as $a^{-6}$. Among all the components commonly considered (e.g. dust: $a^{-3}$, radiation: $a^{-4}$) this is the most singular behavior in the approach to a singularity, while still marginally bearable quantum mechanically. In the presence of anisotropy the wavefunction of the universe is separable into subspaces of given anisotropy charges, much like the partial wave expansion of a scattering problem into angular momentum components. This charge multiplies the most singular part of the effective potential (22), which corresponds to the critical case $\beta=2$. As shown in the last section, in this case, the stability of the system depends on the numerical coefficient $\gamma$, or on the total anisotropy charge $Q$. The fate of our observer thus depends on the amount of anisotropy in its surroundings. However, when the correct units are re-established, the catastrophe seems inevitable. From the previous analysis, the critical value of the anisotropy charge $Q$ is $\alpha^{2}$-suppressed,
@@ Acknowledgements.
Acknowledgments — I acknowledge inspiring exchanges with Jose Beltrán Jiménez, Maurizio Gasperini, Steffen Gielen, Alberto Nicolis, Alessandro Podo, Riccardo Rattazzi, Andrew Tolley and Siméon Vareilles. I am particularly indebted to Mehrdad Mirbabayi and Gabriele Veneziano for important remarks on the draft. This work received support from the French government under the France 2030 investment plan, as part of the Initiative d’Excellence d’Aix-Marseille Université - A*MIDEX (AMX-19-IET-012). It was also supported by the “action thématique” Cosmology-Galaxies (ATCG) of the CNRS/INSU PN Astro.

### v2

@@ Abstract
I investigate spacetime singularities from the point of view of the wavefunction of the universe. In order to extend the classical notion of geodesic incompleteness one has to include the proper time of an observer as a degree of freedom in the Wheeler DeWitt equation. This leads to a Schrödinger equation along the observer worldline. Near the singularity, as in the classical BLK treatment, I ignore spatial gradients and effectively describe the spacetime around the worldline in the mini-superspace approximation. Then the problem proves identical to a spherically symmetric scattering of a quantum particle off a central potential and singularity avoidance is tantamount to unitary evolution for this system. Standard types of matter (dust, radiation) correspond to regular potentials and thus lead to a bounce. The most singular component, spatial anisotropy, is associated to a conserved charge and yields a negative inverse-square potential—like standard angular momentum, but with opposite sign. This potential is critical, in that the unitarity of the evolution depends on the actual numerical factor in front of it, i.e., on the anisotropy charge.
@@ I Introduction
However, this conclusion might be too pessimistic. As long as the gravitational wavefunction maintains its support on smooth sub-Planckian metrics it should be under the control of the low-energy theory, but can still display features that cannot be reproduced by a point-like classical system. For example, quantum tunneling is made possible by the non-vanishing width of the wavefunction, which allows it to “feel” regions of the potential away from its peak. In this case, a weak-coupling semiclassical treatment is indeed available, but it applies to the Euclidean path integral, so it is not directly associated to the classical (Lorentzian) evolution. In gravity, semiclassical techniques apparently oblivious to UV physics have greatly contributed to the understanding of black hole thermodynamics (e.g. Gibbons and Hawking (1977); Hawking and Page (1983); Almheiri et al. (2021) to cite a few).
@@
However, as emphasized above, it should be the proper time of an observer, not some arbitrarily chosen field, that diagnoses a singularity. As the WdW equation only deals with dynamical fields we should introduce the trajectory of the observer and its proper time $\tt{t}$} as new degrees of freedom. While we send the observer into the singularity we can compute $\Psi(q^{a},{\tt{t}})$ along the worldline and obtain a probability conditional to the actual proper time $\tt{t}$} of the observer. Equivalently, we can compute the expectation values of the $q^{a}$ variables at given $\tt{t}$}. This seems the correct starting point to extend the notion of singularity—or avoidance thereof—into the quantum regime. Without incorporating the clock variable $\tt{t}$}, inspection of the “standard” WdW wavefunction $\Psi(q^{a})$ generally leads to inconclusive results.
To clarify this, let me consider a very stupid example with only two fields, the scale factor $a$ and a scalar $\phi$. Say that $\Psi(a,\phi)$ turns out to be peaked along the field space trajectory $\phi=a$. Such a wavefunction has a clear relational interpretation Piazza and Vareilles (2025b). If, e.g., we use $a$ as “time”, we have, approximately, $\langle\phi\rangle=a$. Does this model contain a singularity or a bounce?
Only when $\tt{t}$} is added as a degree of freedom can we address this question. From $\Psi(a,\phi,{\tt{t}})$, one can compute the average values of $a$ and $\phi$ as functions of $\tt{t}$}. The probe should only slightly perturb the system, so $\Psi$ is still be peaked on a trajectory belonging to the two-dimensional plane $a=\phi$. However, this can include very disparate situations. We could have, say, $\langle a\rangle=\langle\phi\rangle={\tt{t}}^{2}+1$, which clearly represents a bounce. Or, we could find that both $\langle a\rangle$ and $\langle\phi\rangle$ tend to infinity at ${\tt{t}}=0$—more precisely, that $\Psi$ does not evolve unitarily in $\tt{t}$}. In this case, we would simply conclude that the low-energy theory alone is not able to resolve the singularity.
@@ II The Schrödinger equation
Here I review how the inclusion of the proper time of the observer into the WdW formalism leads to a Schrödinger equation along the worldline Piazza and Vareilles (2025a). The use of the Schrödinger equation and/or proper time in quantum cosmology has been advocated in a number of papers (e.g. Teitelboim (1983); Unruh and Wald (1989); Peter (2018); Malkiewicz et al. (2022); Burns et al. (2022); Kaya (2023); Bergeron et al. (2025)). Of particular relevance to the present work are Refs. Vachaspati et al. (2007); Greenwood and Stojkovic (2008); Wang et al. (2009); Saini and Stojkovic (2014) for application to singularities.
Quid est Tempus? — Classical GR’s answer to this question is unambiguous, at least locally, i.e. along an observer’s worldline. The invariant line element, or proper time, $d{\tt{t}}$ is obtained by contracting the metric tensor $g_{\mu\nu}$ along the observer’s trajectory—see eq. (2) below. The actual procedures that the observer needs to implement in order to measure proper time are not of concern here. By definition, a good clock must tick homogeneously in the $\tt{t}$} variable.
@@
Quantum mechanically, the variance associated with $\tt{t}$} can be interpreted as the uncertainty to set the origin of our clock variable along the worldline. Or, as Witten nicely puts it Witten (2024),
@@
This gauge corresponds to fixing the spatial coordinates $x^{i}$ in such a way that the shifts $N_{i}=0$. This can be done by identifying a congruence of geodesic observers, among which the observer provided with the clock (2) sits at $x^{i}=0$. The formal way to do so is to introduce a non-relativistic fluid type of matter, described by three scalar fields $\phi^{I}$ where $\phi^{I}=constant$ represent the trajectories of the volume elements of the fluid, or of the observers inside the congruence. By adopting the unitary gauge choice $\phi^{I}=x^{i}$, the momentum constraint then implies $N_{i}=0$ (e.g. Nicolis et al. (2022), App. D). The Hamiltonian constraint is the WdW equation, that I discuss shortly.
@@
where $N{\cal H}$ is the Hamiltonian of the system without the clock. In the standard WdW approach one poses ${\cal H}\Psi(q^{a})=0$ ($q^{a}$ are now mini-superspace quantum variables). By including the clock into the system, the WdW equation becomes ${\cal H}_{\rm tot}\Psi(q^{a},{\tt{t}})=0$. Upon use of the canonical relation $p_{\tt{t}}=-\,i\partial_{\tt{t}}$ this is just the Schrödinger equation as advertized,
@@ III Approaching the singularity
A convenient parameterization — In the absence of anisotropic stresses the metric $h_{ij}$ can be conveniently parametrized as (e.g. Nicolis et al. (2022))
@@
where $\Delta^{\!(2)}=\partial_{+}^{2}+\partial_{-}^{2}$ is the Laplacian in the flat two-dimensional anisotropy space and, as opposed to the schematic (7), all factors of $\alpha$ and $L$ have been made explicit. However, by measuring time in units of $L$, we can effectively set $L=1$ from now on.
@@
As predicted, anisotropies contribute a negative effective potential and act as a centripetal force. This is analogous to having $\ell(\ell+1)<0$ in the Hydrogen atom. I now review what is the maximally singular behavior for the potential to be compatible with unitary evolution. I will show that typical matter ($w=0$ for non-relativistic matter, $w=1/3$ for radiation) contributes to the potential in a regular way and that the only possible source of instability is given by anisotropy, depending on its actual amount, i.e. on the value of $Q$.
@@ IV Unitarity
with $\gamma\geq 0$ and $\beta\geq 0$. Classically, this system is unstable because the Hamiltonian is unbounded from below. This allows potential energy to be converted indefinitely into kinetic energy, leading to runaway singular solutions, as in the case of a radial infall toward the origin. Quantum mechanically, however, the system (23) is regular if Landau and Lifshits (1991)
@@
The RHS of the above inequality is itself a function bounded from below if $\beta<2$, or if $\gamma<1/4$ in the marginal case $\beta=2$. Remarkably, this simple heuristic argument reproduces the correct power-law behavior of the potential and, in $d=3$, also the critical numerical value of $\gamma$ as in (24). According to the classical analysis of Landau and Lifshits (1991) a potential more singular than (24) simply leads to the infall of the quantum particle into the center.
@@
When $\beta>2$ the second term on the LHS of (28) dominates, and it cannot be set to zero or compensated in any obvious way. What happens in this case is that the phase of the wavefunction increases indefinitely while approaching the singularity. This can be seen Perelomov and Popov (1970) already at leading order in the semi-classical expansion $R=e^{i(S_{0}/\hbar+S_{1}+\dots)}$, which gives
@@ V Discussion
Classically, the energy density associated with spatial anisotropy scales as $a^{-6}$. Among all the components commonly considered (e.g. dust: $a^{-3}$, radiation: $a^{-4}$) this is the most singular behavior in the approach to a singularity. However, as shown such a behavior is still marginally bearable quantum mechanically. In the presence of anisotropy the wavefunction of the universe is separable into subspaces of given anisotropy charges, much like the partial wave expansion into angular momentum components. This charge multiplies the most singular part of the effective potential (22), which corresponds to the critical case $\beta=2$. As shown in the last section, in this case, the stability of the system depends on the numerical coefficient $\gamma$, or on the total anisotropy charge $Q$. The fate of our observer thus depends on the amount of anisotropy in its surroundings. However, when the correct units are re-established, the catastrophe seems inevitable. From the previous analysis, the critical value of the anisotropy charge $Q$ is $\alpha^{2}$-suppressed,
@@ Acknowledgements.
Acknowledgments — I acknowledge inspiring exchanges with Jose Beltrán Jiménez, Maurizio Gasperini, Steffen Gielen, Alberto Nicolis, Alessandro Podo, Riccardo Rattazzi, Andrew Tolley and Siméon Vareilles. I am particularly indebted to Mehrdad Mirbabayi and Gabriele Veneziano for important remarks on the draft. This work received support from the French government under the France 2030 investment plan, as part of the Initiative d’Excellence d’Aix-Marseille Université - A*MIDEX (AMX-19-IET-012). It was also supported by the “action thématique” Cosmology-Galaxies (ATCG) of the CNRS/INSU PN Astro and by the Agence Nationale de la Recherche under the grant ANR-24-CE31-6963-01.